OpenStudy (anonymous):

the following equaton has three solutions. look for the GCF then use the quadratic formula to find all solutions. 3x^3+1x^2+1x=0

OpenStudy (anonymous):

x is the gcf so, x (3x^2+x=1)=0 can you do it from here?

OpenStudy (anonymous):

i factored out the x but now i dont know what to do with the quadratic formula

OpenStudy (anonymous):

sorry, there was a typo x (3x^2+x+1)=0

OpenStudy (anonymous):

so i use the 3x^2+x+1 for the quadratic formula?

OpenStudy (anonymous):

look, if x=0, then the equation is true. so that is one solution. Yes then use the qudratic formula to find the other two solutions

OpenStudy (anonymous):

will you be ok now?

OpenStudy (anonymous):

i got \[(-1+\sqrt{11})/2\]

OpenStudy (anonymous):

but its telling that not correct

OpenStudy (anonymous):

hold on while i work it out

OpenStudy (anonymous):

you did the sutbraction wrong under the rad

OpenStudy (anonymous):

and it is -1+ or - the rad/2

OpenStudy (anonymous):

still there?

OpenStudy (anonymous):

yes i dont understand

OpenStudy (anonymous):

-1 +-rad(1-4*3*1)/2*3 -1+-rad (-11)/6